OpenStudy (anonymous):

(Complex Numbers) Find each quotient: 8i/(2+2i)

hartnn (hartnn):

can you find the conjugate of 2+2i first ?

hartnn (hartnn):

because you'll need to multiply the numerator and denominator by the conjugate of 2+2i

OpenStudy (anonymous):

Yeah it's (2-2i), I multiplied the numerator and denominator by it and my end result was: (16+16i)/8, but the answer key in the textbook says it's suppose to be: 2+2i. That's why I know something's not right. Is it just simplified that way because of the 16/8=2?

hartnn (hartnn):

yes, thats correct. when you factor out 8 from the numerator, you get 8(2+2i)/8 = just 2+2i :)

OpenStudy (nurali):

\[\frac{ 8i }{ 2+2i }\] Multiply and Divided by 2-2i \[\frac{ 8i(2-2i) }{ (2+2i)(2-2i) }\] \[\frac{ 16i-16i^2 }{ 4-4i+4i-4i^2 }\] \[i^2=-1\] \[\frac{16i+16 }{ 8 }\] Taking Common \[\frac{ 8(2+2i) }{ 8 }\] \[2+2i\]

OpenStudy (anonymous):

Yeah I see it. Thank you!

OpenStudy (nurali):

My Pleasure.